A Photon and an alpha particle are accelerated through a potential difference of 100V .the ratio of the wavelength associated the proton to that associated with an alpha particle is?

Grade:12th Pass

2 Answers

Vikas TU
11 years ago

K1   = h / underoot 2 x m x 100............1)


 

K2 = h/underoot 2 x 2m x 100..............2)

 

1)/2) = K1/K2 = root2 : 1

 

 

plz approve!

 

DEV MUKHERJEE
11 years ago

 de broglie  wavelength  is given by  

λ= h/sqrt?(2 m k)

hence λi/λ2 =sqrt  (  m alpha k alpha/ m proton  k proton)

=  2sqrt2